Jane took 15 min to drive her boat upstream to water ski. Coming back later, at the same boat speed, took her 5 min. If the current in that part of the river is 10km/hr, what is her boat speed in still water?
Answer by vksarvepalli(154)

let the speed of the boat in still water be v km/hr and the one trip distance be d kms
while upstream the net speed will be (v-10)km/hr
time taken to drive upstream = d/(v-10) hrs i.e., = 15 min = 1/4 hrs
and net speed downstream = (v+10)km/hr
time taken to drive downstream = d/(v+10) hrs = 5 min = 1/12 hrs
so, d/(v-10)=1/4 and d/(v+10)=1/12
dividing both the equations
(v+10)/(v-10) = 3/1
so v+10=3v-30
2v=40
thus v=20
boat speed in still water=20km/hr Ans.